hello it is a word problem .the product of two consecutive positive odd integers is 323. Find the integers.i do not get what i do first. thank you very much.
Answer by user_dude2008(1862)

x(x+2)=323

x^2+2x=323

x^2+2x-323=0

(x+19)(x-17)=0

x+19=0 or x-17=0

x=-19 or x=17

Numbers are: 17, 19